A drug to lower a patient’s blood pressure is ordered to be administered in a liter D5W/0.45%NaCL and administered over 24 hours. If a 143 lb patient is to receive 0.5 mcg/kg/minute, how many milligrams of the drug must be added to this IV infusion? How is this worked?

Answer is 46.8
Answer #1

ANSWER.

* Calculation amount(mg) of drug must be added to this IV infusion.

       Ordered dose = 0.5 mcg/kg/min.

       Weight of the patient = 143 lbs.

Convert lbs to kg

              1 lbs = 0.454 kg

      Weight of the patient = 143 * 0.454

                                        = 64.9 kg.

    Dose required for a patient weight 64.9 kg = 0.5 * 64.9

                                                                        = 32.5 mcg/min.

             Convert mcg to mg for calculation

                       1 mcg = 0.001 mg.

So

     milligrams(mg) of drug must be added to the IV infusion

                                                   for this patient = 32.5 * 0.001

                                                                                       = 0.0325 mg/min

                            (24 hours = 24 * 60 = 1440 min)

         Drug must be added to administer for 24 hours = 0.0325 * 1440 min (24hrs).

                                                                                     = 46.8 mg/day.